AnimeGhoul8863:

Suppose C and D represent two different school populations where C > D and C and D must be greater than 0. Which of the following expressions is the largest? Explain why. Show all work necessary. (C + D)2 2(C + D) C2 + D2 C2 − D2

AnimeGhoul8863:

I am officially Confused to the max e.e

AnimeGhoul8863:

@Shadow

Study2Learn:

I believe it's "C2 + D2"

AnimeGhoul8863:

ok now have to figure out how yew got dat e.e cause i need steps

Study2Learn:

I wanted to elaborate on it however, I'm not certain if I evaluated it correctly.. you should try asking Voca to check this question out. Maybe shadow would know how to do it as well.

AnimeGhoul8863:

@Vocaloid Could you help me for a second e.e

Study2Learn:

@Vocaloid

AnimeGhoul8863:

@dude

Study2Learn:

@dude

AnimeGhoul8863:

what was that?

dude:

Okay we can just try to substitute any values into this Since c is greater than D, we can set C = 2 and D =1 You can test out all options \((C + D)^2\) \((2 + 1)^2\) \(2(C + D)\) \(2(2 + 1) \) \(C^2 + D^2\) \(2^2 + 1^2\) \(C^2 - D^2\) \(2^2 - 1^2\)

AnimeGhoul8863:

ohhhhhhhhhhhh ok let me try

AnimeGhoul8863:

\[(2+1)^2=3^2=9\] \[2(2+1)=2*3=6\] \[2^2+1^2=4+1=5\] \[2^2-1^2=4-1=3\] So the largest Equation is A

dude:

Right

AnimeGhoul8863:

Your a god no one has ever made math sound that simple before

dude:

Haha